MR Bank has the following balance sheet:

 

Table 3: MR Bank balance sheet

Assets

Required Reserves X            

Excess Reserves R50 million



Liabilities

Deposits            Y            



The required reserve ratio is 10%. What is the value of MR bank’s deposits?

 


R 45 million.


R 10 million.


R 55.6 million.


R 50 million.


1
Expert's answer
2021-08-31T16:20:28-0400
"solution"

"required \ reserve \ of \ 10\%\\
excess \ reserve\ of \ 50\ million \\
A\\
excess \ reserve=actual \ reserve\ -required \ reserve\\

deposit=45m+50m=95\\
B\\
50m+10m=60m\\
C\\
50m+55.6m=106.6m\\
D\\
50m+50m=100m"